# Answering the Unanswerable Is to Err Knowingly: Analyzing and Mitigating Abstention Failures in Large Reasoning Models

### v1

@@ 6 Conclusion and future work
In this paper, we investigate the failure of LRMs to abstain from answering unanswerable questions, despite having the cognitive ability to detect the unanswerability. We identify a misalignment between the model’s internal cognition and its external response behavior. We propose a lightweight, two-stage method that significantly improves abstention behavior without harming reasoning performance. Future work aims to explore training-time alignment strategies to improve abstention fidelity.
@@ Appendix B Implementation Details about Experiments
Our method consists of two main components. For Cognitive Monitoring, we train a lightweight linear probe on 2,000 pairs of answerable and unanswerable questions sampled from the SUM dataset (Song, Shi, and Zhao 2025), and validate it on an additional 200 pairs. For the reasoning process of each question, we sample 1,000 token-level activations as input to the probe. The linear probe is trained for 75 epochs with a batch size of 16,384 and a learning rate of 3e-5  (Li et al. 2023). We evaluate the classification performance of the probe using the full reasoning process on both the SUM and UMWP test sets. The results are shown in Table 5, where the F1 score is computed with a threshold of 0.5. The results demonstrate that the probe achieves good classification performance and generalizes well across datasets. In our method, the probe is used to predict the answerability of a question based on the generated reasoning content during the inference process. A question is predicted as unanswerable when the probe’s output exceeds a threshold $t$. We set $t$ = 0.6 for the SUM dataset and $t$ = 0.5 for the UMWP dataset.
LRMs
Best
Layer
SUM UMWP
@@ Appendix C Full Results for Further Discussions
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
@@
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$

### v3

@@ 6 Conclusion and Future Work
We investigate the failure of LRMs to abstain from answering unanswerable questions, despite having the cognitive ability to detect the unanswerability. We identify a misalignment between the model’s internal cognition and its external response behavior. We propose a lightweight, two-stage method that significantly improves abstention behavior without harming reasoning performance. Future work aims to explore training-time alignment strategies to improve abstention fidelity.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Nos. 62272219 and 62406136).
@@ Appendix B Implementation Details about Experiments
Our method consists of two main components. For Cognitive Monitoring, we train a lightweight linear probe on 2,000 pairs of answerable and unanswerable questions sampled from the SUM dataset (Song et al. 2025), and validate it on an additional 200 pairs. For the reasoning process of each question, we sample 1,000 token-level activations as input to the probe. The linear probe is trained for 75 epochs with a batch size of 16,384 and a learning rate of 3e-5  (Li et al. 2023). We evaluate the classification performance of the probe using the full reasoning process on both the SUM and UMWP test sets. The results are shown in Table 5, where the F1 score is computed with a threshold of 0.5. The results demonstrate that the probe achieves good classification performance and generalizes well across datasets. In our method, the probe is used to predict the answerability of a question based on the generated reasoning content during the inference process. A question is predicted as unanswerable when the probe’s output exceeds a threshold $t$. We set $t$ = 0.6 for the SUM dataset and $t$ = 0.5 for the UMWP dataset.
LRMs Best Layer SUM UMWP
@@ Appendix C Full Results for Further Discussions
Correct abstention $\uparrow$ Hallucinated answer $\downarrow$ Cognitive fixation $\downarrow$ Acc $\uparrow$ Correct abstention $\uparrow$ Hallucinated answer $\downarrow$ Cognitive fixation $\downarrow$ Acc $\uparrow$
@@
Correct abstention $\uparrow$ Hallucinated answer $\downarrow$ Cognitive fixation $\downarrow$ Acc $\uparrow$ Correct abstention $\uparrow$ Hallucinated answer $\downarrow$ Cognitive fixation $\downarrow$ Acc $\uparrow$